# Yamaji effect in models of underdoped cuprates

### v1

@@ Appendix C Form factor of inter-layer hopping
Figure 8: Resistivity $\rho_{zz}$ calculated by the ancilla theory with $t_{z}(\mathbf{k})=t_{z}(1+[\cos(k_{x})-\cos(k_{y})]^{2}/4)$, as a function of $\theta$ and $\phi$ for (a) $\omega_{c}\tau=1.3$ and (b) $\omega_{c}\tau=3.0$. (c) $\rho_{zz}(\theta=90^{\circ},\phi)$ as a function of $B$ and $\phi$. Here $B_{0}$ is chosen as the magnetic field such that $\omega_{c}\tau=1.0$. (d) The slope slope $1/a_{1}$ as a function of $\phi$.
Figure 9: Resistivity $\rho_{zz}$ calculated within FL* for LSCO parameters at doping $p=0.21$, with interlayer hopping $t_{z}(\mathbf{k})=t_{z}[\cos(k_{x})-\cos(k_{y})]^{2}/4$, shown as a function of $\theta$ and $\phi$ for (a) $\omega_{c}\tau=0.4$ and (b) $\omega_{c}\tau=1.0$.
Figure 10: Resistivity $\rho_{zz}$ calculated within FL* for LSCO parameters at doping $p=0.21$, with interlayer hopping $t_{z}(\mathbf{k})=t_{z}[\cos(k_{x})-\cos(k_{y})]^{2}/4$, shown as a function of $\theta$ and $\phi$ for $\omega_{c}\tau=0.4$. Panels (a) and (b) correspond to (a) $c_{0}=\cos(\pi/10),c_{\mathbf{k}}+\sin(\pi/10),\psi_{\mathbf{k}}$, and (b) $c_{0}=\cos(\pi/5),c_{\mathbf{k}}+\sin(\pi/5),\psi_{\mathbf{k}}$, respectively.
@@
Appendix D ADMR in LSCO
In this appendix, we show the magnetoresistance curve calculated by ALM in the LSCO sample[4] at doping $p=0.21$, with $a_{\mathrm{lat}}=3.75$$\text{\AA }$ and $c_{\mathrm{lat}}=6.60$$\text{\AA }$. Here we choose $t=0.16$ eV, $t^{\prime}/t=-0.1364$ and $t^{\prime\prime}/t=0.0682$ for the $c$ electron, $t_{\psi}/t=-1$, $t_{\psi}^{\prime}/t=0.1364$ and $t_{\psi}^{\prime\prime}/t=0.0682$ for the $\psi$ electron, and a hybridization $\Phi/t=0.006$. The interlayer hoppings in Eq. 9 are $t_{z,0}=0$ and $t_{z,1}/t=0.0325$. The calculated resistivity is shown in Fig. 9. For smaller $\omega_{c}\tau=0.4$, $\rho_{zz}$ decreases monotonically with $\theta$ and reaches its minimum at $\theta=90^{\circ}$. For larger $\omega_{c}\tau=1.0$, $\rho_{zz}$ first increases to a maximum before decreasing again. The appearance of this maximum signals the onset of a Yamaji-type oscillation, although the first Yamaji angle is smaller than in the isotropic case (with $t_{z,0}$ dominated) due to the strong momentum dependence of $t_{z}(\mathbf{k})$.
@@ Appendix D ADMR in LSCO
Appendix E A brief introduction to the Yamaji effect.
@@
E.1 Chamber’s formula of conductivity
@@
$\delta f(\mathbf{k},t)=-e\mathbf{E}\cdot\mathbf{v}_{\mathrm{ret}}(\mathbf{k}(t),t)\tau\left(-\frac{\partial f}{\partial\epsilon}\right),$ (12)
@@
$\mathbf{v}_{\mathrm{ret}}(\mathbf{k}(t),t)=\frac{1}{\tau}\int_{-\infty}^{t}dt^{\prime}\mathbf{v}(\mathbf{k}(t^{\prime}))e^{-(t-t^{\prime})/\tau}\,.$ (13)
@@
$J_{i}=-2e\int\frac{d^{3}k}{(2\pi)^{3}}f(\mathbf{k},0)v_{i}=2e^{2}E_{j}\int\frac{d^{3}k}{(2\pi)^{3}}\left(-\frac{\partial f}{\partial\epsilon}\right)\int_{-\infty}^{0}dt^{\prime}v_{i}(\mathbf{k}(0))v_{j}(\mathbf{k}(t^{\prime}))e^{t^{\prime}/\tau}\,$, (14)
@@
$\sigma_{ij}=\frac{e^{2}}{4\pi^{3}}\int_{\mathrm{FS}}d^{2}k\mathcal{D}(\mathbf{k})\int_{-\infty}^{0}dt^{\prime}v_{i}(\mathbf{k}(0))v_{j}(\mathbf{k}(t^{\prime}))e^{t^{\prime}/\tau}\,$, (15)
where $\mathcal{D}(\mathbf{k})$ is the local density of states, and $\mathbf{k}(t)$ parametrizes the position along the Fermi-surface orbit. The subscript “FS” indicate integration over the full 3d Fermi surface. In all the calculations in this work, we assume $t_{\perp}$ is much smaller than the in-plane hopping, such that the effect of $v_{z}(B_{x}-B_{y})$ can be neglected. All the cyclotron orbits share the same period in this condition. We also assume $\mathcal{D}(\mathbf{k})\approx\mathcal{D}_{2d}(\mathbf{k})$, with the later fully determined by the in-plane dispersion $\epsilon_{2d}(k_{x},k_{y})$.
E.2 Yamaji effect
@@
where $\epsilon_{3d}(\mathbf{k})$ if the full 3d dispersion and $\epsilon_{2s}(k_{x},k_{y})$ is the in-plane dispersion, $c_{\mathrm{lat}}$ is the lattice constant along the $z$ direction. When $t_{\perp}$ is much smaller than the in-plane hopping, all the cyclotron orbits share the same period, and $\mathcal{D}(\mathbf{k})=\mathcal{D}_{2d}(\mathbf{k})$. The integration over the Fermi surface can replaced by an integration over $k_{z}$ and time $t$. The conductivity along the $z$ direction is then
@@
where ${k}_{\parallel}(t)$ denotes the in-plane component of the momentum along the cyclotron orbit projected to the in plane magnetic field direction $\mathbf{B}_{\parallel}$.

### v3

@@ Appendix C Form factor of inter-layer hopping
Figure 8: Resistivity $\rho_{zz}$ calculated by the ancilla theory with $t_{z}(\mathbf{k})=t_{z}(1+[\cos(k_{x})-\cos(k_{y})]^{2}/4)$, as a function of $\theta$ and $\phi$ for (a) $\omega_{c}\tau=1.3$ and (b) $\omega_{c}\tau=3.0$. (c) Extracted Yamaji angle $\theta_{\mathrm{Yamaji}}$ as a function of the uniform component percentage. (d) The slope $1/a_{1}$ as a function of $\phi$.
Figure 9: Out-of-plane resistivity $\rho_{zz}$ within the ALM framework [(a)-(d)] and SDW framework [(e)-(f)] for various scattering rate anisotropy models. (a) Momentum-space distribution of the scattering rate $\tau_{0}/\tau(\mathbf{k})=0.15+0.5\cos^{2}\phi_{\mathbf{k}}$ for the FL* pocket, where $\phi_{\mathbf{k}}$ denotes the azimuthal angle. The color scale represents the magnitude of $\tau_{0}/\tau(\mathbf{k})$.(b) Angular dependence of the corresponding resistivity $\rho_{zz}$ derived from the scattering profile in (a). (c) Distribution of the scattering form factor $\tau_{0}/\tau(\mathbf{k})=1/[1-0.9\tanh(k_{d}/0.01)]$ for the FL* pocket, where $k_{d}$ is the momentum distance to the folded Brillouin zone boundary. (d) Calculated resistivity $\rho_{zz}$ as a function of $\theta$ for the scattering model defined in (c). (e) Momentum-space distribution of the scattering rate $\tau_{0}/\tau(\mathbf{k})=0.15+0.5\cos^{2}\phi_{\mathbf{k}}$ for the SDW pocket. (f) Angular dependence of the resulting resistivity $\rho_{zz}$ for the SDW pocket case.
Figure 10: Resistivity $\rho_{zz}$ calculated within SDW theory with $\mathbf{Q}=(\pi,\pi,\pi)$, shown as a function of $\theta$ and $\phi$ for (a) $\omega_{c}\tau=0.5$ and (b) $\omega_{c}\tau=1.5$.
Figure 11: Resistivity $\rho_{zz}$ calculated within FL* for LSCO parameters at doping $p=0.21$, with interlayer hopping Eq. (15), shown as a function of $\theta$ and $\phi$ for (a) $\omega_{c}\tau=0.4$ and (b) $\omega_{c}\tau=1.0$.
Figure 12: Resistivity $\rho_{zz}$ calculated within FL* for LSCO parameters at doping $p=0.21$, with interlayer hopping Eq. (15), shown as a function of $\theta$ and $\phi$ for $\omega_{c}\tau=0.4$. Panels (a) and (b) correspond to (a) $c_{0}=\cos(\pi/10)c_{\mathbf{k}}+\sin(\pi/10)\psi_{\mathbf{k}}$, and (b) $c_{0}=\cos(\pi/5)c_{\mathbf{k}}+\sin(\pi/5)\psi_{\mathbf{k}}$, respectively.
@@
Appendix D Form factor of scattering time $\tau$.
In the primary calculations presented in the main text, we assumed a uniform scattering time $\tau$ across the entire Fermi pocket. However, because quasiparticles derive their spectral weight from different components at various segments of the Fermi surface, the scattering rate is expected to exhibit momentum dependence. Specifically, near the corners of the Fermi pockets, quasiparticles undergo strong scattering mediated by the bosonic field $\Phi$ within the ALM framework, or by the order parameter $\Delta$ in the SDW theory.
In this appendix, we examine the robustness of our results by considering two distinct momentum-space distributions for the scattering time $\tau(\mathbf{k})$, as illustrated in Fig. 9. First, we consider an anisotropic scattering rate that reaches its maximum at the pocket corners, defined by:
$\tau_{0}/\tau(\mathbf{k})=0.15+0.5\cos^{2}\phi_{\mathbf{k}}~,$ (13)
where $\phi_{\mathbf{k}}$ denotes the azimuthal angle of $\mathbf{k}$. Second, specifically for the FL* phase, we implement a distribution where the scattering rate is higher for the ancilla fermions compared to the conventional $c$-fermions:
$1/\tau(\mathbf{k})=1/(1-0.9\tanh(k_{d}/0.01))~,$ (14)
where $k_{d}$ represents the momentum-space distance from $\mathbf{k}$ to the folded Brillouin zone boundary. We find that in both models, and for both the FL* and SDW frameworks, the specific distribution of the scattering time $\tau(\mathbf{k})$ has a negligible impact on the overall results. Key features, including the Yamaji angles and the qualitative distinctions between the FL* and SDW phases, remain robust against scattering anisotropy.
Appendix E SDW with $\mathbf{Q}=(\pi,\pi,\pi)$.
In considering the $c$-axis resistivity within the SDW framework, it is important to carefully account for the $c$ component of the SDW folding momentum. While the spin correlation along $c$ is generally antiferromagnetic, it is much weaker than the in-plane correlations. In the main text, we focus on the case with $\mathbf{Q}=(\pi,\pi,0)$. Here, we also examine the alternative possibility that $\mathbf{Q}=(\pi,\pi,\pi)$.
The resulting ADMR curve is shown in Fig. 10, which differs qualitatively from the experimental observation. The ADMR curve decreases as a function of $\theta$ for small $\theta$. This behavior arises from the folding of the interlayer hopping $t_{z}$, which causes the inner and outer halves of the SDW pocket acquire opposite signs of $t_{z}$, effectively introduceing an anisotropy in the interlayer hopping. At higher magnetic fields, a Yamaji peak emerges at a larger angle, but the overall lineshape remains inconsistent with experiment.
Appendix F ADMR in LSCO
In this appendix, we show the magnetoresistance curve calculated by ALM in the LSCO sample[12] at doping $p=0.21$, with $a_{\mathrm{lat}}=3.75$$\text{\AA }$ and $c_{\mathrm{lat}}=6.60$$\text{\AA }$. Here we choose $t=0.16$ eV, $t^{\prime}/t=-0.1364$ and $t^{\prime\prime}/t=0.0682$ for the $c$ electron, $t_{\psi}/t=-1$, $t_{\psi}^{\prime}/t=0.1364$ and $t_{\psi}^{\prime\prime}/t=0.0682$ for the $\psi$ electron, and a hybridization $\Phi/t=0.006$. The interlayer hopping is taken as
$t_{z}(\mathbf{k}_{2d})=\frac{1}{4}t_{z,1}[\cos k_{x}-\cos k_{y}\bigr]^{2}\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}~,$ (15)
where $t_{z,1}/t=0.0325$. To account for the interlayer shift of neighboring $\text{CuO}_{2}$ planes in LSCO, the $t_{z}$ term is further multiplied by a factor of $\cos(k_{x}/2)\cos(k_{y}/2)$. The calculated resistivity is shown in Fig. 11. For smaller $\omega_{c}\tau=0.4$, $\rho_{zz}$ decreases monotonically with $\theta$ and reaches its minimum at $\theta=90^{\circ}$. For larger $\omega_{c}\tau=1.0$, $\rho_{zz}$ first increases to a maximum before decreasing again. The appearance of this maximum signals the onset of a Yamaji-type oscillation, although the first Yamaji angle is smaller than in the isotropic case (with $t_{z,0}$ dominated) due to the strong momentum dependence of $t_{z}(\mathbf{k}_{2d})$.
@@ Appendix F ADMR in LSCO
Appendix G A brief introduction to the Yamaji effect.
@@
G.1 Chamber’s formula of conductivity
@@
$\delta f(\mathbf{k},t)=-e\mathbf{E}\cdot\mathbf{v}_{\mathrm{ret}}(\mathbf{k}(t),t)\tau(\mathbf{k}(t))\left(-\frac{\partial f}{\partial\epsilon}\right),$ (18)
@@
$\mathbf{v}_{\mathrm{ret}}(\mathbf{k}(t),t)=\frac{1}{\tau(\mathbf{k}(t))}\int_{-\infty}^{t}dt^{\prime}\mathbf{v}(\mathbf{k}(t^{\prime}))e^{-\int_{t^{\prime}}^{t}\frac{1}{\tau(\mathbf{k}(t^{\prime\prime}))}dt^{\prime\prime}}\,.$ (19)
@@
$J_{i}=-2e\int\frac{d^{3}k}{(2\pi)^{3}}f(\mathbf{k},0)v_{i}=2e^{2}E_{j}\int\frac{d^{3}k}{(2\pi)^{3}}\left(-\frac{\partial f}{\partial\epsilon}\right)\int_{-\infty}^{0}dt^{\prime}v_{i}(\mathbf{k}(0))v_{j}(\mathbf{k}(t^{\prime}))e^{-\int_{t^{\prime}}^{0}\frac{1}{\tau(\mathbf{k}(t^{\prime\prime}))}dt^{\prime\prime}}\,$, (20)
@@
$\sigma_{ij}=\frac{e^{2}}{4\pi^{3}}\int_{\mathrm{FS}}d^{2}k\mathcal{D}(\mathbf{k})\int_{-\infty}^{0}dt^{\prime}v_{i}(\mathbf{k}(0))v_{j}(\mathbf{k}(t^{\prime}))e^{-\int_{t^{\prime}}^{0}\frac{1}{\tau(\mathbf{k}(t^{\prime\prime}))}dt^{\prime\prime}}\,$, (21)
where $\mathcal{D}(\mathbf{k})$ is the local density of states, and $\mathbf{k}(t)$ parametrizes the position along the Fermi-surface orbit. The subscript “FS” indicate integration over the full 3d Fermi surface. In all the calculations in this work, we assume $t_{\perp}$ is much smaller than the in-plane hopping, such that the effect of $v_{z}(B_{x}-B_{y})$ can be neglected. All the cyclotron orbits share the same period in this condition. We also assume $\mathcal{D}(\mathbf{k})\approx\mathcal{D}_{2d}(\mathbf{k})$, with the later fully determined by the in-plane dispersion $\epsilon_{2d}(k_{x},k_{y})$. In the main text, we ignore the momentum dependence of $\tau(\mathbf{k})$ and approximate $\tau$ as the average scattering time.
G.2 Yamaji effect
@@
where $\epsilon_{3d}(\mathbf{k})$ if the full 3d dispersion and $\epsilon_{2s}(k_{x},k_{y})$ is the in-plane dispersion, $c_{\mathrm{lat}}$ is the lattice constant along the $z$ direction. When $t_{\perp}$ is much smaller than the in-plane hopping, all the cyclotron orbits share the same period, and $\mathcal{D}(\mathbf{k})=\mathcal{D}_{2d}(\mathbf{k}_{2d})$. The integration over the Fermi surface can replaced by an integration over $k_{z}$ and time $t$. The conductivity along the $z$ direction is then
@@
where we assume a constant scattering time $\tau$, $T$ is the period of the quasi-particle cyclotron motion, and ${k}_{\parallel}(t)$ denotes the in-plane component of the momentum along the cyclotron orbit projected to the in plane magnetic field direction $\mathbf{B}_{\parallel}$.